OpenStudy (anonymous):

0 medals for this answerer. (a2b)4∗(ab)2

OpenStudy (anonymous):

Hey! I live in MEdina too, lol... OH right?

OpenStudy (anonymous):

4a8b * 2a2b = 8a^216b^2

OpenStudy (anonymous):

\[(a ^{2}b)^{4}*(ab)^{?}\]

OpenStudy (anonymous):

that question mark should be 2

OpenStudy (anonymous):

Ohh so 4 and 2 are exponents...you did not write the raise.

OpenStudy (anonymous):

yeah, oops... BTW, sorry.... Medina is probably ur last name.

OpenStudy (anonymous):

It's the question, per say, I need to simplify the expression.

OpenStudy (anonymous):

\[(a^2b)^4∗(ab)^2 = answer\]\[ a^8b^4 * a^2b^2 = answer\] \[a^8b^4 * a^2b^2 = a ^{10} b^6\]

OpenStudy (anonymous):

Thank you so much!!!!!!!

OpenStudy (anonymous):

No prob homie

OpenStudy (anonymous):

Btw, can 2m/3n be simplified?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

thx so much.... see ya 'round, got a final 2 study for